1) We have a population of frogs with a mean weight of 0.9 pounds. We select four frogs and find their average weight. What do you suppose it will be?

2)Two events are independent of each other. If the probability of the first is 0.6, and the probability of the second is 0.2, what is the probability that both occur?

3)We are to select a president, vice-president, secretary, and treasurer out of a club with 15 members. Assuming that no one can hold more than one office, how many selections are possible?

4) We have 73 women and 62 men. If we select a person at random, what is the probability that the person is a man?

5) The probability that it will rain on a given day is 0.23. What is the probability that it will not rain on a given day?
Answer #1

1)

It will be 0.9

2)

Probability both occur = 0.6 * 0.2

=.0.12

3)

Number of ways = 15 * 14 * 13 * 12

= 32760

4)

Probability = 62 / 62+73

= 0.4593

5)

P = 1 - 0.23

= 0.77

Please let me know in case of any doubt.

Thanks in advance!


Please upvote!